Compute the cost of not taking the following cash discounts. (Use a 360-day year. Do not round intermediate calculations. Input your final answers as a percent rounded to 2 decimal places.)    Cost of Lost Discount

Cost of Lost Discount
  a. 3/19, net 60. %
  b. 2/19, net 45. %
  c. 4/19, net 45. %
  d. 3/16, net 190. %
Answer #1

Cost of lost discount = Disc. % /(100%-Disc.%) * 360 / ( Due – disc. period)

a)   Cost of lost discount = 3% /(97%) *360/(60-19) = 27.16%

b)   Cost of lost discount = 2% /(98%) *360/(45-19) = 28.26%

c)   Cost of lost discount = 4% /(96%) *360/(45-19) = 57.69%

Cost of lost discount = 3% /(97%) *360/(190-16) =6.40%